LTO Reports
NOAO

KPNO LTO Reports
are available here        

Click here for last night's (161208) :
(default)

Or, enter desired date (yymmdd):
     

                 


LTO Report

Kitt Peak Nightly LTO Report For UT: 2016-12-08

Report Time (MST): 06:33                    For access to all LTO-Reports go to: http://www-kpno.kpno.noao.edu/cgi-bin/ltoreport


4 Meter Weather Averages
From: UT = 01:00 (MST=6:00 PM) To: UT = 13:00 (MST=6:00 AM)
Elements Mean High Low AvDev
Air Temperature (°C)3.54.42.80.3
Dewpoint Temperature (°C)0.53.0-15.31.9
Relative humidity (%)821002311.5
Wind Speed (mph)11.422.31.73.3
Wind Direction (°)203341260.0
4M Dome Temperature (°C)4.77.22.91.1
4M Dome Dewpoint (°C)-2.60.1-7.61.5
4M Primary Mirror (°C)5.15.54.90.1

 
Weather Comments: First half was mostly cloudy with thick cirrus and low clouds moving in at the beginning of the night. After midnight the low clouds moved out and we were left with clear skies. However, humidity was high most of the night (>90%) except on the north end of the mountain where after 00:45 hrs, the relative humidity ranged between 89% and 45%.
Recognized Hours: ( 11.5 )      18 degree twilight + 0.5 hr.

Observer Information
Telescope Observer Prop. # Instrument Detector Seeing
4 MeterMilisavljevic, Crespo, Negro 2016B-0901 KOSMOS E2V 1.0
WIYNColon, Barentsen 2016B-0196 WHIRC N/A N/A
2.1 MeterSzafruga 2016B-2007 N/A N/A N/A
FeedNo Observer --- N/A N/A N/A
0.9 MeterGraykowski, Hui 2016B-0070 HDI N/A N/A

Time Usage During The Night
Telescope Hours Used Night's Activity   Hours Lost Time Lost
Due To:
Time of
Night (MST)
4Meter
5.5 Scientific Research   4.75 Weather 20:00 - 06:00
1.25 Checkout        
0.25 Extra Science Time        
WIYN
1.5 Scientific Research 10.0 Weather 20:00 - 06:00
2Meter
11.5 T and E        
Feed
0.0 Scientific Research      
W0_9m
1.5 Other   10.0 Weather 20:00 - 06:00

Observing Status
4 MeterLost most of the night to high humidity and thick cirrus. Opened up for about 30 minutes at the beginning of the night to verify pointing and for the observer to check out the instrument on sky. After midnight the humidity decreased and we were able to open again.Lost about 10 minutes around 4am when the MPK GUIs all dissapeared -A quick restart fixed the issue. Seeing measured between 1 and 1.7 arc-seconds, with a typical value of 1.
WIYNObserved out of focus so no seeing measured. There was an issue with the moss rotator hitting the hard limit - see MSR#25747. No dark time was lost since it happened while closed due to the high humidity.
2.1 MeterT&E for the recent TCS upgrade. No on-sky work.
FeedNot Scheduled
0.9 MeterNo answer when called at 5:45 am.

Comments4meter was the only lucky telescope to observe. The high humidity prevented the other scopes to get on sky.

Reported by:Christian Soto



Weather Plots